OpenStudy (anonymous):

Given the points P = (2; 5) and Q = (-3; 2) answer each of the parts (a)-(d). (a) Find an equation of the line passing through P and Q. (b) Find the midpoint of the segment connecting P and Q. (c) Find an equation of the right bisector of the segment passing through P and Q. (d)Find an equation of the line through (1;

OpenStudy (anonymous):

Easy!!

OpenStudy (anonymous):

a. 3x-5y+19=0 b. (-0.5,3.5) c. 5x+3y-8=0 d. 3x-5y+7=0

OpenStudy (anonymous):

Nice